40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 244     Entry time: Thu Oct 22 14:33:50 2020
Author: Marie K. 
Type: Progress 
Category: BS BRDs 
Subject: BS BRDs for LHO are assembled and tuned 

The 5 BS BRDs for installation at LHO are now assembled and tuned in the LLO Optics Lab. The target frequencies are 17.79 Hz and 26.06 Hz (ref alog 49643).

The masses I had to use for the Bounce mode are slightly heavier than I was expecting (above 10.10g instead of 9.90g). The masses for the Roll mode are in target. I deliberately tuned the BRDs to lower frequencies (between -1% and -2% below the target) to anticipate the drift that shifts the frequencies to higher values. After 2 days, the drifts seem lower than was previoulsy observed at CIT (about 0.1% per day compared to 0.5% per day) with some of the BRDs exhibiting no drift at all. That might be good news? Monitoring will continue over the next weeks.

The Q values fluctuate from measurement to measurement, but they are essentially close to 100.

I am using the LLO B&K vibrometer kit, with an older version of the software that Stuart help me to setup, to measure the resonant frequencies. The air flow from the flow bench is currently off.

Attachment 1: LHO_BS_BRDs.jpg  1.046 MB  Uploaded Thu Oct 22 15:53:11 2020  | Hide | Hide all
LHO_BS_BRDs.jpg
Attachment 2: Test_setup.jpg  1.772 MB  Uploaded Thu Oct 22 15:53:23 2020  | Hide | Hide all
Test_setup.jpg
Attachment 3: Monitoring_2020_10_22.png  778 kB  Uploaded Thu Oct 22 15:57:56 2020  | Hide | Hide all
Monitoring_2020_10_22.png
ELOG V3.1.3-